OpenStudy (anonymous):

In a right triangle, find the length of the side not given. ___ b=1, c=sqr root of 26

OpenStudy (saifoo.khan):

5

OpenStudy (saifoo.khan):

Use, sqrt(26)^2 = 1^2 + a^2 26 - 1 = a^2 a^2 = 25 a = 5

OpenStudy (anonymous):

Thank you